# Local expression of fractional corner charges in obstructed atomic insulators and its application to vertex-transitive polyhedra with arbitrary genus

### v1

@@ V.1 Review of fractional disclination charges in two dimensions
Here we review the fractional disclination charges in two dimensions based on Ref. [PhysRevB.101.115115]. First, we briefly explain the concept of disclinations. A two-dimensional $C_{n}$ symmetric crystal can be divided into $n$ equivalent sectors, each of which has the internal angle $\frac{2\pi}{n}$ at the center of the crystal, and is mutually related by the $C_{n}$ symmetry. By adding or removing such sectors into/from the crystal, a zero-dimensional defect appears at the center of the crystal, which is called a disclination.
@@
Consider the $C_{4}$-symmetric square lattice with disclination of $\Omega=-\frac{\pi}{2}$ as an example. In this case, there are two types of disclinations, i.e., $\bm{a}]^{(4)}=0$ in Fig. 7(a) and $\bm{a}]^{(4)}=1$ in Fig. 7(b), in which the sum $a_{1}+a_{2}$ from $\bm{a}$ over the closed path is even or odd, respectively. Namely, the holonomy in Fig. 7(a) is given by $\hat{r}(\frac{\pi}{2})\hat{t}_{2\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{2\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{2\bm{e}_{1}}=\hat{t}_{-2\bm{e}_{1}}\hat{r}(-\frac{\pi}{2})$, leading to $\bm{a}]^{(4)}=0$, where we used $\hat{r}(R)\hat{t}_{\bm{a}}=\hat{t}_{R\bm{a}}\hat{r}(R)$, while the holonomy in Fig. 7(b) is given by $\hat{r}(\frac{\pi}{2})\hat{t}_{\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{\bm{e}_{1}}=\hat{t}_{-\bm{e}_{1}}\hat{r}(-\frac{\pi}{2})$, leading to $\bm{a}]^{(4)}=1$. In the same way, $\bm{a}]^{(2)}$ for the Frank angle $\Omega=\pi$ takes values in $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, corresponding to the cases with $a_{1}$ and $a_{2}$ being odd or even. Likewise, $\bm{a}]^{(3)}$ for the Frank angle $\Omega=\pm\frac{2\pi}{3}$ takes values in $\mathbb{Z}_{3}$, i.e., $a_{1}-a_{2}\ (\text{mod}\ 3)$, and $\bm{a}]^{(6)}$ for the Frank angle $\Omega=\pm\frac{2\pi}{6}$ is zero.
Figure 8: Wyckoff positions in (a) $C_{2}$-, (b) $C_{4}$-, (c) $C_{3}$-, and (d) $C_{6}$-symmetric systems. The points with the same colors belong to the same Wyckoff positions. $\bm{a}_{1}\ (=a\bm{e}_{1})$ and $\bm{a}_{2}\ (=a\bm{e}_{2})$ are primitive lattice vectors of lattice constant $a$. The gray regions are unit cells.
@@
Now we review the fractional disclination charges in two dimensions [PhysRevB.101.115115]. Consider two-dimensional $C_{n}$-symmetric systems, and let $\Delta\omega\ (=n_{\omega}-m_{\omega})$ denote the total charge of the ionic and electronic charges at the WP $\omega$ in the unit of |e|-|e| as discussed in Sec. IIA. In this case, the bulk polarizations $\bm{P}^{(n)}\ (n=2,4,3,6)$ (per elementary charge $|\absolutevalue{e}$) of the $C_{n}$-symmetric systems, whose unit cells are shown in Fig. 8, are given by
@@
Then, the fractional disclination charges $Q^{(n)}_{\text{dis}}\ (n=2,4,3,6)$ in $C_{n}$-symmetric lattices with disclination of $\Omega$ (per elementary charge $|\absolutevalue{e}$) are given by [PhysRevB.101.115115]
@@ VI Coupling constant in Wen-Zee action
In the previous section, we show that the disclination charge is universally given by Eq. (24) by constructing the relationship between the two-dimensional disclination charges and the corner charges in the corresponding VTSPs. Meanwhile, two-dimensional disclination charges can be described by the Wen-Zee action [PhysRevB.107.195153, PhysRevB.106.L241113]. Here we derive the value of the coupling constant in the Wen-Zee action by considering the disclination charge given by Eq. (24).
First, we briefly review the Wen-Zee action. Let $M$ be a $(2+1)$-dimensional spacetime manifold: $M=\Gamma\times\mathbb{R}$, where $\Gamma$ is a two-dimensional spatial manifold and $\mathbb{R}$ is the time. In this case, the Wen-Zee action represents the coupling between the gauge field $A\ (=A_{\mu}dx^{\mu})\ (\mu=0,1,2)$ and the background curvature $R$ in the manifold $\Gamma$ in the context of the quantum Hall effect [PhysRevLett.69.953] as follows:
@@ VII Conclusion
Moreover, we show that the disclination charges in the two-dimensional systems are universally given by Eq. (24). They equal to the corner charges in the corresponding VTSPs if they share the same degree of sharpness. In particular, we also find that for the disclination patterns with the Frank angle $\Omega\ (>0)$, the corresponding polyhedra has higher genus ($g>0$).

### v2

@@ V.1 Review of fractional disclination charges in two dimensions
Here we review the fractional disclination charges in two dimensions based on Ref. [27]. First, we briefly explain the concept of disclinations. A two-dimensional $C_{n}$ symmetric crystal can be divided into $n$ equivalent sectors, each of which has the internal angle $\frac{2\pi}{n}$ at the center of the crystal, and is mutually related by the $C_{n}$ symmetry. By adding or removing such sectors into/from the crystal, a zero-dimensional defect appears at the center of the crystal, which is called a disclination.
@@
Consider the $C_{4}$-symmetric square lattice with disclination of $\Omega=-\frac{\pi}{2}$ as an example. In this case, there are two types of disclinations, i.e., $\bm{a}]^{(4)}=0$ in Fig. 7(a) and $\bm{a}]^{(4)}=1$ in Fig. 7(b), in which the sum of $a_{1}+a_{2}$ from $\bm{a}=a_{1}\bm{e}_{1}+a_{2}\bm{e}_{2}$, taken over the vectors forming the closed path, is even or odd, respectively. Namely, the holonomy in Fig. 7(a) is given by $\hat{r}(\frac{\pi}{2})\hat{t}_{2\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{2\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{2\bm{e}_{1}}=\hat{t}_{-2\bm{e}_{1}}\hat{r}(-\frac{\pi}{2})$, leading to $\bm{a}]^{(4)}=0$, where we used $\hat{r}(R)\hat{t}_{\bm{a}}=\hat{t}_{R\bm{a}}\hat{r}(R)$, while the holonomy in Fig. 7(b) is given by $\hat{r}(\frac{\pi}{2})\hat{t}_{\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{\bm{e}_{1}}\hat{r}(\frac{\pi}{2})\hat{t}_{\bm{e}_{1}}=\hat{t}_{-\bm{e}_{1}}\hat{r}(-\frac{\pi}{2})$, leading to $\bm{a}]^{(4)}=1$. In the same way, $\bm{a}]^{(2)}$ for the Frank angle $\Omega=\pi$ takes values in $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, corresponding to the cases with $a_{1}$ and $a_{2}$ being odd or even. Likewise, $\bm{a}]^{(3)}$ for the Frank angle $\Omega=\pm\frac{2\pi}{3}$ takes values in $\mathbb{Z}_{3}$, i.e., $a_{1}-a_{2}\ (\text{mod}\ 3)$, and $\bm{a}]^{(6)}$ for the Frank angle $\Omega=\pm\frac{2\pi}{6}$ is zero.
Figure 8: WPs in (a) $C_{2}$-, (b) $C_{4}$-, (c) $C_{3}$-, and (d) $C_{6}$-symmetric systems. The points with the same colors belong to the same WPs. $\bm{a}_{1}\ (=a\bm{e}_{1})$ and $\bm{a}_{2}\ (=a\bm{e}_{2})$ are primitive lattice vectors of lattice constant $a$. The gray regions are unit cells.
@@
Now we review the fractional disclination charges in two dimensions [27]. Consider two-dimensional $C_{n}$-symmetric systems, and let $\Delta\omega\ (=n_{\omega}-m_{\omega})$ denote the total charge of the ionic and electronic charges at the WP $\omega$ in the unit of |e|-|e| as discussed in Sec. IIA. In this case, the bulk polarizations $\bm{P}^{(n)}\ (n=2,4,3,6)$ (per elementary charge $|\absolutevalue{e}$) of the $C_{n}$-symmetric systems, whose unit cells are shown in Fig. 8, are given by [27]
@@
Then, the fractional disclination charges $Q^{(n)}_{\text{dis}}\ (n=2,4,3,6)$ in $C_{n}$-symmetric lattices with disclination of $\Omega$ (per elementary charge $|\absolutevalue{e}$) are given by [27]
@@ VI Coupling constant in Wen-Zee action
In the previous section, we show that the disclination charge is universally given by Eq. (24) by constructing the relationship between the two-dimensional disclination charges and the corner charges in the corresponding VTSPs. Meanwhile, two-dimensional disclination charges can be described by the Wen-Zee action [31, 28]. Here we derive the value of the coupling constant in the Wen-Zee action by considering the disclination charge given by Eq. (24).
First, we briefly review the Wen-Zee action. Let $M$ be a $(2+1)$-dimensional spacetime manifold: $M=\Gamma\times\mathbb{R}$, where $\Gamma$ is a two-dimensional spatial manifold and $\mathbb{R}$ is the time. In this case, the Wen-Zee action represents the coupling between the gauge field $A\ (=A_{\mu}dx^{\mu})\ (\mu=0,1,2)$ and the background curvature $R$ in the manifold $\Gamma$ in the context of the quantum Hall effect [54] as follows:
@@ VII Conclusion
[…]
Moreover, we show that the disclination charges in the two-dimensional systems are universally given by Eq. (24). They equal to the corner charges in the corresponding VTSPs if they share the same degree of sharpness. In particular, we also find that for the disclination patterns with the Frank angle $\Omega\ (>0)$, the corresponding polyhedra has higher genus ($g>0$). We note that the corresponding VTSPs are determined only by the degree of sharpness for the disclination center. In that case, all the vertices in the VTSP should be connected by certain three-dimensional point group symmetry due to the vertex transitiveness of the VTSP.